# The Cosmological Impact of Brane-Chern-Simons Massive Gravity

### v1

@@ Abstract
In this paper, we present a novel extension of massive gravity theory; the Brane-Chern-Simons massive gravity theory. We explore the cosmological implications of this theory by deriving the background equations and demonstrating the existence of self-accelerating solutions. Interestingly, our theory suggests the existence of self-accelerating mechanisms that originate from an effective cosmological constant, leading to intriguing possibilities for understanding the nature of cosmic acceleration. Furthermore, we perform a tensor perturbation analysis to investigate the propagation of gravitational waves in this framework. We derive the dispersion relation for gravitational waves and study their behavior in the Friedmann-Lemaître-Robertson-Walker cosmology within the context of Brane-Chern-Simons massive gravity. Utilizing the latest Union2 type Ia supernovae dataset comprising 557 SNIa events, we provide observational support for our theoretical framework, indicating that the Brane-Chern-Simons massive gravity theory is consistent with cosmological observations.
@@ 1 Introduction
The main objectives of this work are to demonstrate the existence of a self-accelerating solution in Brane-Chern-Simons massive gravity which is related to the effective cosmological constant and to perform a perturbation analysis. In particular, we will derive the modified dispersion relation for gravitational waves within this theory. Furthermore, to constrain the parameters of the theory, we employ observational data from Type Ia supernovae. By utilizing Bayesian statistical methods, specifically focusing on minimizing $X^{2}$, we can rigorously compare the predictions of the theory with these observational datasets.
The paper is structured as follows: In Section 2, we introduce the Brane-Chern-Simons massive gravity framework, derive the background equations of motion, and extract self-accelerating solutions. Section 3, delves into the cosmological perturbation analysis, focusing on tensor perturbations and the dispersion relation of gravitational waves. Section 4, involves testing the solution of Brane-Chern-Simons massive gravity against the latest Union2 type Ia supernovae (SNIa) dataset. Finally, in Section 5, we summarize our key findings and discuss future research directions.
@@ 2 Brane-Chern-Simons Massive Gravity
$S=\frac{M_{\rm Pl}^{2}}{2}\int d^{4}x\Bigg{\{}\sqrt{-g}\bigg{[}R+\varepsilon RV(\varphi)-2\kappa V(\varphi)+2{m}_{g}^{2}U(\mathcal{K})\bigg{]}\Bigg{\}},$
also, $m_{g}$ is the mass of graviton. In the following, we introduce the $U({\cal K})$ of this action separately. It is clear that the mass of the graviton comes up with the potential $U$ which consists of three parts.
@@
$\mathcal{K}^{\mu}_{\nu}=\delta^{\mu}_{\nu}-\sqrt{g^{\mu\alpha}f_{\alpha\nu}},$ (4)
where $f_{\alpha\nu}$ is the fiducial metric, which is defined through
$f_{\alpha\nu}=\partial_{\alpha}\phi^{c}\partial_{\nu}\phi^{d}\eta_{cd}.$ (5)
Here $g^{\mu\nu}$ is the physical metric, $\eta_{cd}$ is the Minkowski metric with $c,d=0,1,2,3$ and $\phi^{c}$ are the Stueckelberg fields which are introduced to restore general covariance. According to our cosmological application purpose, we adopt the Friedman-Lemaître-Robertson-Walker (FLRW) Universe. So, the general expression of the corresponding dynamical and fiducial metrics are given as follows,
$g_{\mu\nu}$ $={\rm diag}\left[-N^{2},a^{2},a^{2},a^{2}\right],$ (6)
$f_{\mu\nu}$ $={\rm diag}\left[-\dot{f}(t)^{2},1,1,1\right].$ (7)
Here it is worth pointing out that $N$ is the lapse function of the dynamical metric, and it is similar to a gauge function. Also, it is clear that the scale factor is represented by $a$, and $\dot{a}$ is the derivative with respect to time. Furthermore, the lapse function relates the coordinate-time $dt$ and the proper-time $d\tau$ via $d\tau=Ndt$ Scheel:1994yr ; Christodoulakis:2013xha . Function $f(t)$ is the Stueckelberg scalar function whereas $\phi^{0}=f(t)$ and $\frac{\partial\phi^{0}}{\partial t}=\dot{f}(t)$ Arkani-Hamed:2002bjr . Therefore, the point-like Lagrangian of the Brane-Chern-Simons massive gravity in FLRW cosmology is
$\mathcal{L}=-\frac{3M_{\rm Pl}^{2}\big{[}1+\varepsilon V(\varphi)\big{]}\dot{a}^{2}}{N^{2}a^{2}}-\kappa M_{\rm Pl}^{2}V(\varphi)+\frac{m_{g}^{2}(a-1)M_{\rm Pl}^{2}}{a^{3}N}\Bigg{\{}\bigg{(}\alpha_{3}+\alpha_{4}+a\big{[}a(\alpha_{4}+4\alpha_{3}$
$+6)-2\alpha_{4}-5\alpha_{3}-3\big{]}\bigg{)}N-\bigg{(}a\big{[}a(\alpha_{4}+3\alpha_{3}+3)-2\alpha_{4}-3\alpha_{3}\big{]}+\alpha_{4}\bigg{)}\dot{f}(t)\Bigg{\}}.$
In order to simplify expressions later, we define
@@ 2.1 Background Equations of Motion
In order to achieve a constraint equation we should take the unitary gauge into consideration, which means that we choose $f(t)=t$. The significance of the unitary gauge lies in the fact that on the classical level, the unphysical fields could be eliminated from the Lagrangian with the use of gauge transformations Grosse-Knetter:1992tbp .
In this procedure, a constraint equation can be derived by varying with respect to $f$. So, the equation is given by
$\frac{\delta\mathcal{L}}{\delta f}=3m_{g}^{2}M_{\rm Pl}^{2}\dot{a}\bigg{\{}\alpha_{3}+\alpha_{4}-2a(1+2\alpha_{3}+\alpha_{4})+a^{2}(3+3\alpha_{3}+\alpha_{4})\bigg{\}}=0.$ (10)
@@
$\frac{1}{M_{\rm Pl}^{2}a^{3}}\frac{\delta\mathcal{L}}{\delta N}=3H^{2}(1+\varepsilon V(\varphi))-\kappa V(\varphi)+\frac{m_{g}^{2}(a-1)}{a^{3}}\bigg{[}\alpha_{3}+\alpha_{4}+a\big{(}-3$
$-5\alpha_{3}-2\alpha_{4}+a(6+4\alpha_{3}+\alpha_{4})\big{)}\bigg{]}=0.$ (11)
@@
$\frac{1}{3M_{\rm Pl}^{2}a^{2}N}\frac{\delta\mathcal{L}}{\delta a}=3H^{2}\big{[}1+\varepsilon V(\varphi)\big{]}-\kappa V(\varphi)+\frac{2\dot{H}(1+\varepsilon V(\varphi))}{N}+\frac{m_{g}}{r^{2}N^{3}}\Bigg{\{}N\bigg{[}(1+2r)\big{(}1$
$+2\alpha_{3}+\alpha_{4}\big{)}+rN\big{(}rN(6+4\alpha_{3}+\alpha_{4})-(2+r)(3+3\alpha_{3}+\alpha_{4})\big{)}\bigg{]}-\alpha_{3}-\alpha_{4}\Bigg{\}}=0,$
@@
$\frac{\delta S}{\delta\sigma}\dot{\sigma}+\frac{\delta S}{\delta f}\dot{f}-N\frac{d}{dt}\frac{\delta S}{\delta N}+\dot{a}\frac{\delta S}{\delta a}=0.$ (14)
@@ 2.2 Self-Accelerating Background Solutions
In this step, we want to discuss solutions. It could be started with the Stueckelberg constraint in Eq. (10), so, we have
$\alpha_{3}+\alpha_{4}-2a(1+2\alpha_{3}+\alpha_{4})+a^{2}(3+3\alpha_{3}+\alpha_{4})=0,$
it would be suitable to mention that the constant solutions of $a$ lead to the effective energy density and behave similarly to a cosmological constant. As a result, the two solutions of Eq. (2.2) are
$a_{\rm SA}^{\pm}=\frac{1+2\alpha_{3}\pm\sqrt{1+\alpha_{3}+\alpha_{3}^{2}-\alpha_{4}}+\alpha_{4}}{3+3\alpha_{3}+\alpha_{4}}.$ (16)
@@
$3H^{2}\bigg{(}1+\varepsilon V(\varphi)\bigg{)}=\Lambda_{\rm SA}^{\pm}+\kappa V(\varphi).$ (17)
Considering self-accelerating solutions, a condition on the parameter $\varepsilon V(\varphi)$ is provided by the Friedman equation (17). So, we need to consider $\varepsilon V(\varphi)>-1$ to keep the left-hand side of Eq. (17) positive. The importance of this the issue lies in the fact that when we add ordinary matters to the right-hand side, throughout the matter-dominated era, we will have the standard cosmology. It is worth mentioning that the effective cosmological constant from the mass term is
$\Lambda_{\rm SA}^{\pm}\equiv\frac{m_{g}^{2}(1-a_{\rm SA}^{\pm})}{a_{\rm SA}^{3\pm}}\bigg{[}\alpha_{3}+\alpha_{4}+a_{\rm SA}^{\pm}\big{(}-3-5\alpha_{3}-2\alpha_{4}+a_{\rm SA}^{\pm}(6+4\alpha_{3}+\alpha_{4})\big{)}\bigg{]}.$ (18)
According to Eq. (16), the above equation can be written as
$\Lambda_{\rm SA}^{\pm}=\frac{m_{g}^{2}}{(\alpha_{3}+\alpha_{4})^{2}}\Bigg{\{}2\sqrt{1+\alpha_{3}+\alpha_{3}^{2}-\alpha_{4}}\bigg{(}\pm 1\mp\alpha_{4}\pm\alpha_{3}\pm\alpha_{3}^{2}\bigg{)}+3\alpha_{4}(1+\alpha_{3})$
$-3\alpha_{3}(\alpha_{3}+1)-2\alpha_{3}^{3}-2\Bigg{\}}.$ (19)
Therefore, from Eq. (2.1) we have,
$r_{\rm SA}=\frac{m_{g}^{2}N\bigg{[}a_{\rm SA}^{2\pm}\big{(}\alpha_{3}+2\big{)}+N\big{(}2\alpha_{3}+3\big{)}-2a_{\rm SA}^{\pm}N\big{(}\alpha_{3}+2\big{)}-\alpha_{3}-1\bigg{]}}{\big{(}a_{\rm SA}^{\pm}-1\big{)}^{2}}\pm\Bigg{\{}\frac{m_{g}^{2}N^{2}}{\big{(}a_{\rm SA}^{\pm}-1\big{)}^{2}}\bigg{[}m_{g}^{2}\big{(}\alpha_{3}$
$+1\big{)}^{2}+\bigg{(}m_{g}^{2}\big{(}\alpha_{3}(\alpha_{3}+3)+3\big{)}+\kappa V(\varphi)\big{(}\alpha_{3}+1\big{)}-3H^{2}(\alpha_{3}+1)\big{(}\varepsilon V(\varphi)+1\big{)}\bigg{)}N^{2}-N\big{(}\alpha_{3}$
$+1\big{)}\bigg{(}m_{g}^{2}\big{(}2\alpha_{3}+3\big{)}+2\dot{H}\big{(}\varepsilon V(\varphi)+1\big{)}\bigg{)}+a_{\rm SA}^{2\pm}\bigg{(}m_{g}^{2}\big{(}\alpha_{3}+2\big{)}^{2}+\big{[}m_{g}^{2}(\alpha_{3}+3)-\kappa V(\varphi)$
$+3H^{2}(\varepsilon V(\varphi)+1)\big{]}\big{(}N(\alpha_{3}+2)-2\alpha_{3}-3\big{)}N-6\dot{H}+2\big{[}\varepsilon V(\varphi)\big{(}(\alpha_{3}+2)N-2\alpha_{3}-3\big{)}$
$+N(\alpha_{3}+2)-2\alpha_{3}\big{]}\dot{H}\bigg{)}+2a_{\rm SA}^{\pm}\bigg{(}(\alpha_{3}+1)\big{[}m_{g}^{2}\big{(}2\alpha_{3}+5\big{)}-\kappa V(\varphi)+3H^{2}\big{(}\varepsilon V(\varphi)+1\big{)}\big{]}N$
$-m_{g}^{2}N^{2}\big{(}\alpha_{3}+2\big{)}^{2}+(\alpha_{3}+1)\big{(}2\dot{H}(\varepsilon V(\varphi)+1)-m_{g}^{2}(\alpha_{3}+2)\big{)}\bigg{)}\bigg{]}\Bigg{\}}^{\frac{1}{2}}$
×$\times\Bigg{\{}\frac{m_{g}^{2}N^{2}\bigg{[}2a_{\rm SA}^{\pm}\big{(}\alpha_{3}+2\big{)}-2\alpha_{3}-3\bigg{]}}{\big{(}a_{\rm SA}^{\pm}-1\big{)}^{2}}+\bigg{(}\frac{m_{g}^{2}\big{[}3(\alpha_{3}+2)-2(2\alpha_{3}+5)a_{\rm SA}^{\pm}+(\alpha_{3}+3)a_{\rm SA}^{2\pm}\big{]}}{\big{(}a_{\rm SA}^{\pm}-1\big{)}^{2}}$
$-\kappa V(\varphi)+3H^{2}\big{(}\varepsilon V(\varphi)+1\big{)}\bigg{)}N^{3}+2N^{2}\dot{H}\big{(}\varepsilon V(\varphi)+1\big{)}\Bigg{\}}^{-1}.$
Actually, we have used the Stuckelberg equation (2.2) in order to eliminate $\alpha_{4}$.
@@
A key finding of this subsection is the identification of self-accelerating solutions within the theory, characterized by an effective cosmological constant, $\Lambda_{\rm SA}^{\pm}$. Importantly, these solutions do not exhibit strong coupling issues, and they provide a well-behaved description of the accelerated expansion of the universe. This suggests that the Brane-Chern-Simons massive gravity theory offers a viable framework for explaining the late-time cosmic acceleration without encountering the challenges typically associated with strong coupling regimes.

### v2

@@ Abstract
In this paper, we present a novel extension of massive gravity theory; the Brane-Chern-Simons massive gravity theory. We explore the cosmological implications of this theory by deriving the background equations and demonstrating the existence of self-accelerating solutions. Interestingly, our theory suggests the existence of self-accelerating mechanisms that originate from an effective cosmological constant, leading to intriguing possibilities for understanding the nature of cosmic acceleration. Furthermore, we perform a tensor perturbation analysis to investigate the propagation of gravitational waves in this framework. We derive the dispersion relation for gravitational waves and study their behavior in the Friedmann-Lemaître-Robertson-Walker cosmology within the context of Brane-Chern-Simons massive gravity.
@@ 1 Introduction
The main objectives of this work are to demonstrate the existence of a self-accelerating solution in Brane-Chern-Simons massive gravity which is related to the effective cosmological constant and to perform a perturbation analysis. In particular, we will derive the modified dispersion relation for gravitational waves within this theory.
The paper is structured as follows: In Section 2, we introduce the Brane-Chern-Simons massive gravity framework, derive the background equations of motion, and extract self-accelerating solutions. Section 3, delves into the cosmological perturbation analysis, focusing on tensor perturbations and the dispersion relation of gravitational waves. Finally, in Section 4, we summarize our key findings and discuss future research directions.
@@ 2 Brane-Chern-Simons Massive Gravity
$S=\frac{M_{\rm Pl}^{2}}{2}\int d^{4}x\Bigg{\{}\sqrt{-g}\bigg{[}R+\varepsilon RV(\varphi)-2\kappa V(\varphi)+2{m}_{g}^{2}U(\mathcal{K})\bigg{]}\Bigg{\}}+\int d^{4}x\sqrt{-g}\mathcal{L}_{m},$
also, $m_{g}$ is the mass of graviton and $\mathcal{L}_{m}$ is the matter Lagrangian. In the following, we introduce the $U({\cal K})$ of this action separately. It is clear that the mass of the graviton comes up with the potential $U$ which consists of three parts.
@@
$\mathcal{K}^{\mu}_{\nu}=\delta^{\mu}_{\nu}-\sqrt{g^{\mu\alpha}\partial_{\alpha}\phi^{c}\partial_{\nu}\phi^{d}\eta_{cd}},$ (4)
where $g^{\mu\nu}$ is the physical metric, $\eta_{cd}$ is the Minkowski metric with $c,d=0,1,2,3$ and $\phi^{c}$ are the Stueckelberg fields which are introduced to restore general covariance. According to our cosmological application purpose, we adopt the general form of Friedman-Lemaître-Robertson-Walker (FLRW) Universe. So, the general expression of the corresponding dynamical metric is given as follows,
$g_{\mu\nu}dx^{\mu}dx^{\nu}$ $=-N^{2}dt^{2}+a^{2}\Omega_{i,j}dx^{i}dx^{j},$ (5)
$\Omega_{i,j}dx^{i}dx^{j}$ $=dx^{2}+dy^{2}+dz^{2}-\frac{|K|(xdx+ydy+zdz)^{2}}{1+|K|(x^{2}+y^{2}+z^{2})}.$ (6)
Here it is worth pointing out that $N$ is the lapse function of the dynamical metric, and it is similar to a gauge function. Also, it is clear that the scale factor is represented by $a$, and $\dot{a}$ is the derivative with respect to time. Meanwhile, it should be mentioned that $x^{0}=t$, $x^{1}=x$, $x^{2}=y$, $x^{3}=z$; $\mu,\nu=0,...,3$; and $i,j=1,2,3$. Furthermore, $|K|$ is the Gaussian curvature of the space and, the lapse function relates the coordinate-time $dt$ and the proper-time $d\tau$ via $d\tau=Ndt$ Scheel:1994yr ; Christodoulakis:2013xha .
Concerning the scalar fields $\phi^{c},(c=0,...,3)$, the following ansatz are considered,
$\phi^{0}$ $=f(t)\sqrt{1+|K|(x^{2}+y^{2}+z^{2})},$
$\phi^{1}$ $=\sqrt{|K|}f(t)x,$
$\phi^{2}$ $=\sqrt{|K|}f(t)y,$
$\phi^{3}$ $=\sqrt{|K|}f(t)z.$ (7)
where function $f(t)$ is the Stueckelberg scalar function whereas $\frac{\partial f(t)}{\partial t}=\dot{f}(t)$ Arkani-Hamed:2002bjr .
Therefore, the point-like Lagrangian of the Brane-Chern-Simons massive gravity is
$\mathcal{L}=-\frac{aM_{\rm Pl}^{2}}{N}\Bigg{\{}\bigg{(}3|K|+\big{(}3|K|\varepsilon+\kappa a^{2}\big{)}V(\varphi)\bigg{)}N^{2}+3\big{(}1+\varepsilon V(\varphi)\big{)}\dot{a}^{2}\Bigg{\}}+m_{g}^{2}\big{(}a$
$-\sqrt{|K|}f(t)\big{)}M_{\rm Pl}^{2}\Bigg{\{}|K|f(t)^{2}\bigg{(}(\alpha_{3}+\alpha_{4})N-\alpha_{4}\dot{f}(t)\bigg{)}+a^{2}\bigg{(}(6+4\alpha_{3}+\alpha_{4})N$
$-(3+3\alpha_{3}+\alpha_{4})\dot{f}(t)\bigg{)}+\sqrt{|K|}af(t)\bigg{(}(3\alpha_{3}+2\alpha_{4})\dot{f}(t)-(3+5\alpha_{3}+2\alpha_{4})N\bigg{)}\Bigg{\}}+\mathcal{L}_{m}.$
We can assume the matter sector $\mathcal{L}_{m}$ to consist of a perfect fluid with the energy density $\rho_{m}$ and pressure $p_{m}$. In order to simplify expressions later, we define
@@ 2.1 Background Equations of Motion
Without loss of generality, we can assume that $\dot{f}\geq 0$, $f\geq 0$, $a>0$, and $N>0$, at least within the vicinity of the time of interest. Consequently, the equation obtained by varying with respect to $f(t)$ encapsulates all the nontrivial information. In this procedure, a constraint equation can be derived by varying with respect to $f$. So, the equation is given by
$\frac{\delta\mathcal{L}}{\delta f}=3m_{g}^{2}M_{\rm Pl}^{2}\big{(}\dot{a}-\sqrt{|K|}N\big{)}\bigg{\{}a^{2}\big{(}3+3\alpha_{3}+\alpha_{4}\big{)}-2\sqrt{|K|}af\big{(}1+2\alpha_{3}+\alpha_{4}\big{)}$
$+|K|f^{2}\big{(}\alpha_{3}+\alpha_{4}\big{)}\bigg{\}}=0.$ (10)
@@
$\frac{1}{M_{\rm Pl}^{2}a^{3}}\frac{\delta\mathcal{L}}{\delta N}=3H^{2}(1+\varepsilon V(\varphi))-\frac{3|K|(1+\varepsilon V(\varphi))}{a^{2}}-\kappa V(\varphi)+\frac{m_{g}^{2}}{a^{3}}\bigg{[}a^{3}\big{(}\alpha_{4}+4\alpha_{3}+6\big{)}$
$-3a^{2}\sqrt{|K|}f\big{(}\alpha_{4}+3\alpha_{3}+3\big{)}+3af^{2}|K|\big{(}\alpha_{4}+2\alpha_{3}+1\big{)}-|K|^{\frac{3}{2}}f^{3}\big{(}\alpha_{4}+\alpha_{3}\big{)}\bigg{]}=\rho_{m}.$
@@
$\frac{1}{3M_{\rm Pl}^{2}a^{2}N}\frac{\delta\mathcal{L}}{\delta a}=3H^{2}\big{[}1+\varepsilon V(\varphi)\big{]}-\kappa V(\varphi)+\frac{2\dot{H}(1+\varepsilon V(\varphi))}{N}-\frac{|K|}{r^{2}N^{2}}\big{(}1+\varepsilon V(\varphi)\big{)}$
$+\frac{m_{g}^{2}}{r^{2}N^{3}}\Bigg{\{}2\sqrt{|K|}rfN\bigg{(}1+2\alpha_{3}+\alpha_{4}-\big{(}3+3\alpha_{3}+\alpha_{4}\big{)}N\bigg{)}+|K|f^{2}\bigg{(}\big{(}1+2\alpha_{3}$
$+\alpha_{4}\big{)}N-\alpha_{3}-\alpha_{4}\bigg{)}+r^{2}N^{2}\bigg{(}N\big{(}\alpha_{4}+4\alpha_{3}+6\big{)}-\big{(}\alpha_{4}+3\alpha_{3}+3\big{)}\bigg{)}\Bigg{\}}=p_{m},$
@@
$\frac{\delta S}{\delta\varphi}\dot{\varphi}+\frac{\delta S}{\delta f}\dot{f}-N\frac{d}{dt}\frac{\delta S}{\delta N}+\dot{a}\frac{\delta S}{\delta a}=0.$ (14)
@@ 2.2 Self-Accelerating Background Solutions
In this step, we want to discuss solutions. It could be started with the Stueckelberg constraint in Eq. (2.1), so, we have
$\big{(}\dot{a}-\sqrt{|K|}N\big{)}\bigg{\{}a^{2}\big{(}3+3\alpha_{3}+\alpha_{4}\big{)}-2\sqrt{|K|}af\big{(}1+2\alpha_{3}+\alpha_{4}\big{)}+|K|f^{2}\big{(}\alpha_{3}+\alpha_{4}\big{)}\bigg{\}}=0,$
Notice that the above equation has three solutions. The first solution, $\dot{a}=|K|N$, suggests that the physical metric $g_{\mu\nu}$ represents Minkowski spacetime in the open FRW chart. As such, it does not accurately depict the characteristics of our Universe and is thus not a viable solution for describing the cosmos we inhabit Gumrukcuoglu:2011ew . Thus, we eliminate this particular solution, yielding the following expression:
$a^{2}\big{(}3+3\alpha_{3}+\alpha_{4}\big{)}-2\sqrt{|K|}af\big{(}1+2\alpha_{3}+\alpha_{4}\big{)}+|K|f^{2}\big{(}\alpha_{3}+\alpha_{4}\big{)}=0,$ (16)
As a result, the two solutions of Eq. (16) are
$f=\frac{a}{\sqrt{|K|}}\varpi_{\pm},\qquad\varpi_{\pm}=\frac{(1+2\alpha_{3}+\alpha_{4})\pm\sqrt{1+\alpha_{3}+\alpha_{3}^{2}-\alpha_{4}}}{(\alpha_{3}+\alpha_{4})}$ (17)
It is important to note that these two solutions are invalid when $K$ is set to zero. This aligns with the understanding that there exists no non-trivial flat FRW solution under such conditions.
@@
$3H^{2}\bigg{(}1+\varepsilon V(\varphi)\bigg{)}-\frac{3|K|\big{(}1+\varepsilon V(\varphi)\big{)}}{a^{2}}=\rho_{g}+\rho_{m}+\kappa V(\varphi).$ (18)
It is worth mentioning that the effective energy density from the graviton mass term is
$\rho_{g}=-\frac{m_{g}^{2}}{a^{3}}\bigg{[}a^{3}\big{(}\alpha_{4}+4\alpha_{3}+6\big{)}-3a^{2}\sqrt{|K|}f\big{(}\alpha_{4}+3\alpha_{3}+3\big{)}+3af^{2}|K|\big{(}\alpha_{4}$
$+2\alpha_{3}+1\big{)}-|K|^{\frac{3}{2}}f^{3}\big{(}\alpha_{4}+\alpha_{3}\big{)}\bigg{]}.$ (19)
According to Eq. (17), the above equation can be written
$\rho_{g}=\frac{m_{g}^{2}}{\big{(}\alpha_{3}+\alpha_{4}\big{)}^{2}}\bigg{\{}-(1+\alpha_{3})\big{(}2+\alpha_{3}+2\alpha_{3}^{2}-3\alpha_{4}\big{)}\pm\frac{2\big{(}|K|(1+\alpha_{3}+\alpha_{3}^{2}-\alpha_{4})a^{2}\big{)}^{\frac{3}{2}}}{|K|^{\frac{3}{2}}a^{3}}\bigg{\}}.$
By employing Eq. (18), we can express the dynamical equation Eq. (2.1) in an alternative form as follows:
$-\frac{2|K|\big{(}1+\varepsilon V(\varphi)\big{)}}{a^{2}}-\frac{2\dot{H}(1+\varepsilon V(\varphi))}{N}=(\rho_{m}+p_{m})+(\rho_{g}+p_{g}),$ (21)
$p_{g}=\frac{m_{g}^{2}}{r^{2}N^{3}}\Bigg{\{}2\sqrt{|K|}rfN\bigg{(}1+2\alpha_{3}+\alpha_{4}-\big{(}3+3\alpha_{3}+\alpha_{4}\big{)}N\bigg{)}+|K|f^{2}\bigg{(}\big{(}1+2\alpha_{3}$
$+\alpha_{4}\big{)}N-\alpha_{3}-\alpha_{4}\bigg{)}+r^{2}N^{2}\bigg{(}N\big{(}\alpha_{4}+4\alpha_{3}+6\big{)}-\big{(}\alpha_{4}+3\alpha_{3}+3\big{)}\bigg{)}\Bigg{\}}.$ (22)
where $p_{g}$ represents the effective pressure contribution of the graviton mass terms. The equation (21) does not introduce any new insights or information, as it is a direct consequence of the Bianchi identities and the conservation of matter.
Also, we can represent the effective pressure term and the effective density term as shown below:
$p_{eff}=\frac{m_{g}^{2}(1+\alpha_{3})\big{(}2+\alpha_{3}+2\alpha_{3}^{2}-3\alpha_{4}\big{)}}{(\alpha_{3}+\alpha_{4})^{2}}\pm\bigg{[}2m_{g}^{2}\big{(}|K|(1+\alpha_{3}+\alpha_{3}^{2}$
$-\alpha^{4})a^{2}\big{)}^{3/2}|K|^{-3/2}(\alpha_{3}+\alpha_{4})^{-2}\mp a\big{(}|K|+|K|\varepsilon V(\varphi)\big{)}\bigg{]}\times a^{-3}-\kappa V(\varphi)+p_{m},$ (23)
$\rho_{eff}=\frac{m_{g}^{2}}{(\alpha_{3}+\alpha_{4})^{2}}\bigg{[}-(1+\alpha_{3})\big{(}2+\alpha_{3}+2\alpha_{3}^{2}-3\alpha_{4}\big{)}\pm 2\big{(}|K|(1+\alpha_{3}+\alpha_{3}^{2}$
$-\alpha_{4})a^{2}\big{)}^{3/2}|K|^{-3/2}a^{-3}\bigg{]}+\rho_{m},$ (24)
$By\quad considering\quad V(\varphi)=-\frac{|K|}{|K|\varepsilon+2\kappa a^{2}}\Longrightarrow\qquad p_{eff}=-\rho_{eff}.$ (25)
By considering the solutions of the constraint equation (17) and equation (21), the Friedman equation (18) could be rewritten as,
$3H^{2}\bigg{(}1+\varepsilon V(\varphi)\bigg{)}-\frac{3|K|\big{(}1+\varepsilon V(\varphi)\big{)}}{a^{2}}=\tilde{\rho}_{g}+\kappa V(\varphi)+\rho_{m},$ (26)
where
$\tilde{\rho}_{g}=m_{g}^{2}\Lambda_{\pm},$ (27)
$\Lambda_{\pm}\equiv-\frac{1}{(\alpha_{3}+\alpha_{4})^{2}}\bigg{\{}1+\alpha_{3}\pm\sqrt{1+\alpha_{3}+\alpha_{3}^{2}-\alpha_{4}}\bigg{\}}\times\bigg{\{}1+\alpha_{3}^{2}-2\alpha_{4}$
±$\pm(1+\alpha_{3})\sqrt{1+\alpha_{3}+\alpha_{3}^{2}-\alpha_{4}}\bigg{\}}.$ (28)
Equation (26) is analogous to the Friedmann equation describing an open Universe influenced by arbitrary matter and an effective cosmological constant $\tilde{\rho}_{g}=m_{g}^{2}\Lambda_{\pm}$. This cosmological constant arises from the massive graviton term, contributing to the dynamics of the Universe’s expansion.
@@
A key finding of this subsection is the identification of self-accelerating solutions within the theory, characterized by an effective cosmological constant, $m_{g}^{2}\Lambda_{\pm}$. Importantly, these solutions do not exhibit strong coupling issues, and they provide a well-behaved description of the accelerated expansion of the Universe. This suggests that the Brane-Chern-Simons massive gravity theory offers a viable framework for explaining the late-time cosmic acceleration without encountering the challenges typically associated with strong coupling regimes. In the absence of the massive gravity component, i.e., when $m_{g}=0$, the Friedmann equation reduces to the equation of state of standard general relativity, incorporating a scalar potential term, and a possible curvature term. Furthermore, when $m_{g}=0$, $\varepsilon=0$ and $\kappa=0$, the theory recovers the standard general relativity framework. On the background level, our model is equivalent to the standard Lambda-Cold Dark Matter (LCDM) model, which is known to provide an adequate fit to the observational data. Notably, a similar analysis of LCDM model parameter constraints using the Supernova Ia dataset has been previously reported by the Supernova Cosmology Project SupernovaCosmologyProject:2008ojh .